# Flux-balance laws for spinning bodies under the gravitational self-force

### v1

@@ Abstract
The motion of an extended, but still weakly gravitating body in general relativity can often be determined by a set of conserved quantities. If they are sufficient in number, this allows for a complete solution to the equations of motion, much like for geodesic motion. Under the gravitational self-force (relaxing the “weakly gravitating” assumption), these “conserved quantities” evolve with time. This evolution can be calculated using the (local) self-force on the body, but such an approach is computationally intensive. To avoid this, one often instead uses flux-balance laws: relationships between the average evolution (capturing the dissipative dynamics) and the values of the field far away from the body, which are far easier to compute. In the absence of spin, such a flux-balance law has been proven in [1] for any of the conserved action variables appearing in a Hamiltonian formulation of geodesic motion in the Kerr spacetime. In this paper, we attempt to extend this flux-balance law to linear order in spin, in the context of an arbitrary black hole spacetime, and in a way that directly relates the average rates of change to the flux of a conserved current through the horizon and to infinity. In the absence of spin, we reproduce results consistent with those in [1]. To linear order in spin we show that most, but not all, constants of motion at linear order in spin appear to possess flux-balance laws.
@@ I Introduction
While the self-force is often neglected, there exist astrophysical regimes where the gravitational self-force is quite relevant. One such regime is in an extreme mass-ratio inspiral (EMRI): a stellar-mass compact object (which we call the secondary, of mass $m$) orbiting a supermassive black hole (the primary, of mass $M\gg m$). The mass ratio, $\varepsilon\equiv m/M$, measures how strongly the secondary affects the surrounding spacetime, and is therefore a measure of the strength of the gravitational self-force. Due to their low-frequency nature, gravitational waves emitted by these systems are expected to be an important source for the upcoming space-based detector LISA [7, 8, 9]. However, in addition to the ever-present noise, there will be many different sources that all contribute to the signal seen by LISA [7]. The computation of detailed, efficiently-produced gravitational wave models for EMRI signals will be important for their detection, parameter estimation, and any tests of general relativity that they may provide.
@@
It is with this goal in mind that, in recent years, there has been much progress in understanding the gravitational self-force; see for example [10, 11] for reviews of the more pragmatic aspects and progress on the computation of waveforms, or more foundational reviews in [12, 13, 14]. While much of the work has focused on the simplest case, where the secondary has vanishing spin and higher multipoles, astrophysically relevant EMRIs may have secondaries with non-negligible spin. Failure to capture these effects could result in biases in parameter estimation, or in spurious “violations” of general relativity, and so there have been significant recent efforts to compute the effects of spin on self-forced motion (see, for example, [15, 16, 17, 18, 19, 20, 21]).
@@
For black holes, the spin is constrained to be smaller than the square of the mass, and so the effects of spin are second-order in the mass ratio, and for bodies with $O(1)$ compactness, this will also be the case (this is why the spin appears at second order in the discussion of [14]). Because one would also need to generically include the effect of higher spin-induced multipoles (as in Newtonian theory; see for example Sec. 2.4.5 of [22]), we truncate at linear order in spin, and neglect all higher multipoles, which are similarly highly suppressed for compact bodies. Despite the fact that it is somewhat inconsistent, in this paper we neglect the full effects of second-order self force, leave such a discussion to future work, and consider the effects of spin to be of part of the first-order self force.
@@
Up to certain caveats laid out in Footnote 5 below, the motion in this case is described by a sort of generalized equivalence principle: the motion of a self-gravitating extended body is the same as that of a test extended body in an “effective” metric $\breve{g}_{ab}$. This effective metric is constructed from the background metric $g_{ab}$ and a regularized metric perturbation at the location of the body. While this is certainly a simple prescription, solving for the motion in this effective metric is computationally intensive, due to difficulties in constructing this regularized metric perturbation, or indeed any metric perturbation, at the location of the body (see, for example, the extensive discussion in [11]).
@@
An alternative, easier approach, using flux-balance laws, has often been considered throughout the literature. Here, instead of solving for this effective metric near the body to get the exact motion of the body, a flux-balance law relates the average motion of the body to the emitted radiation that reaches the boundaries of the system: infinity and the horizon of the black hole. An example of a flux-balance law appearing in physics folklore is in the decay of the classical atom: since accelerating charges radiate, the emission of energy for an orbiting electron should decrease its orbital energy, resulting in a plunge towards the nucleus.
@@
This argument appeals to a sort of global energy balance, which highlights a key feature of flux-balance laws: they allow for the direct computation not of (even averaged) position or velocity, but of the slowly-evolving, no-longer-conserved quantities of the background motion. In the Kerr spacetime, for example, there are four constants of motion for a geodesic: the square of the mass $m^{2}$, the energy $E$, the azimuthal angular momentum $L_{z}$, and the “Carter constant” $K$ [23]. The Carter constant, which like $m^{2}$ is quadratic in the four-momentum, is related to the square of the orbital angular momentum $L^{2}$ in the limit where the primary black hole has no spin. These are sufficient to determine geodesic motion, but when the body has spin, in addition to generalizations of each of these constants [24, 25], there exists a fifth constant of motion, the Rüdiger constant $Y$ [24]. For any sort of flux balance law to be useful, it must therefore provide average rates of change of these quantities.
@@
Two general approaches have been used in the literature in order to derive flux-balance laws. The first is to use the properties of Green’s functions in the Kerr spacetime to derive relationships between the rates of change of quasi-conserved quantities to the amplitudes of solutions to the Teukolsky equation, which can be written as terms defined at the horizon and infinity. This approach was pioneered in [26] for computations of $E$ and $L_{z}$ without spin, which was then extended to the linear-in-spin case (much later) in [27] (whose derivation was later corrected in [15]). In order to determine the full motion without spin, the evolution of the Carter constant $K$ was computed (once again in terms of asymptotic quantities) in [28, 29], which was later extended in [1] to the case of arbitrary constants of motion for geodesic motion in Kerr. The conserved quantities considered in [1] were the action variables in a Hamiltonian formulation of geodesic motion, which they then show can be used to compute the evolution of $K$.
@@
Recently, however, it was shown in [33] that, at least in the context of a scalar field theory and in the absence of spin (where the evolution of the Carter constant was derived by [34]), one can write the average evolution of the action variables in terms of fluxes of an appropriate conserved current through a local worldtube around the body. This presented a partial physical explanation for the results of [1]. Moreover, by naïvely guessing how to extend this flux-balance law all the way to the horizon and null infinity, a derivation was presented in [33] for a flux balance law, for scalar fields, that was exactly analogous to the results in [1].
This paper begins where [33] ends, and provides an extension of this local flux-balance law beyond the limited scope of that paper. As such, its goals are threefold:
@@
$\left\langle\delta\dot{\Gamma}^{A}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=\Pi^{AB}\left[\mathcal{F}_{B}+\nabla_{B}\langle\mathcal{H}(\tau^{\prime},X)\rangle_{\tau^{\prime}}\right],$ (1)
Without describing the notation explicitly, we will briefly describe this equation’s general features: first, the left-hand side represents a notion of an average rate of change of quantities defined on the worldline, and is defined in Eqs. (22) and (23). For the right-hand side, the first term is a flux of a conserved current through null infinity and the horizon that is defined in terms of metric perturbations in Eq. (110) below, and the second term is a derivative of an average “coincidence” Hamiltonian [defined in Eq. (92)] which generically appears, and is important in the case of resonances [1].
Unfortunately, because of the presence of the coincidence Hamiltonian, this equation does not constitute a flux-balance law relating rates of change on the worldline to asymptotic fluxes of a conserved current, as this term still involves computing the regularized metric perturbation, from which the effective metric $\breve{g}_{ab}$ is constructed, on the worldline. Moreover, the left-hand side is not directly related to averaged rates of change of perturbed conserved quantities. As described in more detail in Sec. II.1 below, it is a covariantly-defined average of the rate of change of $\delta\Gamma^{A}$, the first-order deviation of the self-forced motion from the background trajectory $\Gamma$ through phase space. However, in order to make this average covariantly defined on phase space [like the remaining terms in Eq. (1)], one needs to include a bitensor introduced in [33] called the Hamilton propagator, which significantly complicates the interpretation in the case of spin, as we discuss in Sec. V.2 below.
@@
assuming that the background motion is non-resonant, where $\vartheta^{\alpha}$ represents the angle variables. Unlike Eq. (1), this is a true flux-balance law, as the right-hand side only involves the computation of the metric perturbation $h_{ab}$ at null infinity and the horizon. Moreover, it is practical, as the left-hand side is the average rate of change of the perturbed conserved quantity $\delta P_{\alpha}$, in the usual sense. In particular, this equation holds for non-spinning motion in the Kerr spacetime, and is equivalent to the results of [1]. However, in the case of spin, there no longer exist action-angle variables on phase space, in a sense which we describe in more detail in Sec. V.2 below. As such, our results are much more limited: while we can still recover flux-balance laws for generalizations of $m^{2}$, $E$, $L_{z}$, and $K$, the evolution of the Rüdiger constant $Y$ is less well-behaved.
@@
Following [33], for tensor fields on the phase space manifold $\mathcal{M}$, we use capital Latin letters from the beginning of the alphabet ($A$, $B$, etc.) to denote abstract indices, and capital Hebrew letters from the beginning of the alphabet ($\aleph$, $\beth$, etc.) for coordinate indices. We will also use unhatted Greek letters for indices associated with groupings of four coordinates on phase space, such as the angle variables $\vartheta^{\alpha}$ or the conserved quantities $P_{\alpha}$. Finally, for quantities appearing on the spacetime manifold $M$, we typically denote corresponding quantities on the phase space manifold $\mathcal{M}$ with capitalized versions of the same symbol; for example, the worldline $\gamma$ is determined by a trajectory $\Gamma$ through phase space.
@@ II.1 General framework
We denote the parameters of $\gamma$ and $\Gamma$ by $\tau$, as they will represent the proper time of the body, so that $\dot{\gamma}^{a}$ will be normalized.
@@
$\{f,\{g,h\}\}+\{g,\{h,f\}\}+\{h,\{f,g\}\}=0,$ (5)
or equivalently that the Schouten(-Nijenhuis) bracket of $\Pi^{AB}$ with itself vanishes (see Theorem 10.6.2 of [37]). The equations of motion for $\gamma$ are then given by a Hamiltonian formulation if there exists a function $H$ on $\mathcal{M}$ that satisfies Hamilton’s equations:
@@
Now, we consider the more general case of a pseudo-Hamiltonian system. Here, the idea is that what one would typically think of as a Hamiltonian is not simply a function of a point on phase space $X$, but a function also of a trajectory through phase space $\bar{\Gamma}$ which is determined by a point $\bar{X}$ through which $\bar{\Gamma}$ passes. Following [38, 39, 40], we call such a function a pseudo-Hamiltonian. In the pseudo-Hamiltonian generalization of Hamilton’s equations, one takes a derivative with respect to $X$ at $\bar{X}$ fixed, and then take $\bar{X}$ and $X$ to lie along the same trajectory $\Gamma$ in phase space. These systems allow for dissipation, unlike “true” Hamiltonian systems [41]. [^2] Note that, while they sound somewhat exotic, the idea of pseudo-Hamiltonian systems is far from new (see the discussion and references in [41]). For an example from the gravitational radiation literature, the Burke-Thorne potential for the leading-order post-Newtonian radiation-reaction force (see the discussion in Chapter 12 of [22]), $\Phi^{RR}(x)=\frac{1}{5}x_{i}x_{j}\frac{\mathrm{d}^{5}I^{\langle ij\rangle}}{\mathrm{d}t^{5}},$ can be thought of as the potential energy part of a pseudo-Hamiltonian: the quadrupole moment $I^{ij}$ depends on the trajectory of the object undergoing radiation reaction, which is kept fixed when $\Phi^{RR}$ is differentiated with respect to $x^{i}$ to find the force {to see this, compare Eqs. (12.198) and (12.204) of [22]}.
@@
for some $\tau$, and some trajectory $\Gamma$ in $\mathcal{M}$ that passes through $X$ and obeys Hamilton’s equations for the background Hamiltonian $H$. We then define
@@
when the small parameter $\varepsilon=0$, the system becomes the Hamiltonian for the background motion that determines $\Upsilon$ [^3] In order to describe self-force in the so-called “self-consistent formalism” [42, 14], one would need to instead consider the Hamiltonian flow map associated with the exact motion in the definition of the pseudo-Hamiltonian, which we would denote $\Upsilon_{\Delta}(\varepsilon)$. However, since we are computing the first-order self-force, the self-consistent formalism is not necessary, and so we just use the background Hamiltonian flow map.
@@
$\left.\delta\Gamma^{A}\nabla_{A}f\right|_{\Gamma(\tau)}\equiv\left.\frac{\partial f[\Gamma(\tau;\varepsilon)]}{\partial\epsilon}\right|_{\varepsilon=0}.$ (12)
@@
As shown in Section II.B.2 of [33], the following evolution equation for $\delta\Gamma^{A}$ follows from Eq. (11), using the notation of the pseudo-Hamiltonian formalism (which was only implicit in [33]):
@@
That is, the Poisson bracket structure is fixed, as a function of $\varepsilon$. Note that this is a “gauge-dependent” statement: it only holds in a certain class of coordinates [for example, those in Eq. (43)], and if one performs an $\varepsilon$-dependent coordinate transformation (depending, for example, on the metric perturbation), then Eq. (14) will necessarily change. It is for this reason that we work in coordinates related to those in Eq. (43) by field-independent coordinate transformations.
@@
$\nabla_{A}[f\circ\Upsilon_{\tau^{\prime}-\tau}]=\Upsilon^{A^{\prime}}{}_{A}\nabla_{A^{\prime}}f.$ (16)
In other words, the Hamilton propagator is the pushforward of the Hamiltonian flow map $\Upsilon_{\Delta}$, and is a bitensor that is a one-form at $\Gamma(\tau)$ and a vector at $\Gamma(\tau+\Delta)$. [^4] Note that this definition is slightly different from that originally presented in [33]: there, $\Upsilon^{A^{\prime}}{}_{A}$ was defined such that it was a bitensor at any two points $X$ and $X^{\prime}$ along a given trajectory $\Gamma$, by using a function which gave the difference $\Delta$ in proper times between these two points. This allows one to take derivatives of $\Upsilon^{A^{\prime}}{}_{A}$ with respect to either of $X$ and $X^{\prime}$, entirely independently. However, as we discovered while writing the present paper, such a feature is not required for the derivation of flux-balance laws, and so we do not include this part of the definition.
@@
Finally, we note that, in both Eqs. (23) and (25), care must be taken that the quantity being averaged does, in fact, have a finite and well-defined average. As we will see below in Sec. V.1, this is not always the case, for particular components of these equations; however, in such cases, it will be true that both the left- and right-hand sides of Eq. (23) will both be ill-defined.

### v4

@@ Abstract
The motion of an extended, but still weakly gravitating body in general relativity can often be determined by a set of conserved quantities. Much like for geodesic motion, a sufficient number of conserved quantities allows the motion to be solved by quadrature. Under the gravitational self-force (relaxing the “weakly gravitating” assumption), the motion can then be described in terms of the evolution these “conserved quantities”. This evolution can be calculated using the (local) self-force on the body, but such an approach is computationally intensive. To avoid this, one often uses flux-balance laws: relationships between the average evolution (capturing the dissipative dynamics) and the values of the field far away from the body, which are far easier to compute. In the absence of spin, such a flux-balance law has been proven in [1] for any of the conserved action variables appearing in a Hamiltonian formulation of geodesic motion in the Kerr spacetime. In this paper, we derive a corresponding flux-balance law, to linear order in spin, directly relating average rates of change to the flux of a conserved current through the horizon and out to infinity. In the absence of spin, this reproduces results consistent with those in [1]. To linear order in spin, we construct flux-balance laws for four of the five constants of motion for spinning bodies in the Kerr spacetime, although not in a practical form. However, this result provides a promising path towards deriving the flux-balance law for the (generalized) Carter constant.
@@ I Introduction
While the self-force is often neglected, there exist astrophysical regimes where the gravitational self-force is quite relevant. One such regime is in an extreme mass-ratio inspiral (EMRI): a stellar-mass compact object (which we call the secondary, of mass $m$) orbiting a supermassive black hole (the primary, of mass $M\gg m$). The mass ratio, $\varepsilon\equiv m/M$, measures how strongly the secondary affects the surrounding spacetime, and is therefore a measure of the strength of the gravitational self-force. Due to their low-frequency nature, gravitational waves emitted by these systems are expected to be an important source for the upcoming space-based detector LISA [7, 8, 9, 10]. However, in addition to the ever-present noise, there will be many different sources that all contribute to the signal seen by LISA [7, 10]. The computation of detailed, efficiently-produced gravitational wave models for EMRI signals will be important for their detection, parameter estimation, and any tests of general relativity that they may provide.
@@
It is with this goal in mind that, in recent years, there has been much progress in understanding the gravitational self-force; see for example [11, 12] for reviews of the more pragmatic aspects and progress on the computation of waveforms, or more foundational reviews in [13, 14, 15]. While much of the work has focused on the simplest case, where the secondary has vanishing spin and higher multipoles, astrophysically relevant EMRIs may have secondaries with non-negligible spin. Failure to capture these effects could result in biases in parameter estimation, or in spurious “violations” of general relativity, and so there have been significant recent efforts to compute the effects of spin on self-forced motion (see, for example, [16, 17, 18, 19, 20, 21, 22, 23]).
@@
For black holes, the spin is constrained to be smaller than the square of the mass, and so the effects of spin are second-order in the mass ratio, and for bodies with $O(1)$ compactness, this will also be the case (this is why the spin appears at second order in the discussion of [15]). Because one would also need to generically include the effect of higher spin-induced multipoles (as in Newtonian theory; see for example Sec. 2.4.5 of [24]), we truncate at linear order in spin, and neglect all higher multipoles, which are similarly highly suppressed for compact bodies. Despite the fact that it is somewhat inconsistent, in this paper we neglect the full effects of second-order self-force, leave such a discussion to future work, and consider the effects of spin to be of part of the first-order self-force.
@@
Up to certain caveats laid out in Footnote 9 below, the motion in this case is described by a sort of generalized equivalence principle: the motion of a self-gravitating extended body is the same as that of a test extended body in an “effective” metric $\breve{g}_{ab}$. This effective metric is constructed from the background metric $g_{ab}$ and a regularized metric perturbation at the location of the body. While this is certainly a simple prescription, solving for the motion in this effective metric is computationally intensive, due to difficulties in constructing this regularized metric perturbation, or indeed any metric perturbation, at the location of the body (see, for example, the extensive discussion in [12]).
@@
A more tractable approach, using flux-balance laws, has often been considered throughout the literature. Here, instead of solving for this effective metric near the body to get the exact motion of the body, a flux-balance law relates the average motion of the body to the emitted radiation that reaches the boundaries of the system: infinity and the horizon of the black hole. An example of a flux-balance law appearing in physics folklore is in the decay of the classical atom: since accelerating charges radiate, the emission of energy by an orbiting electron should decrease its orbital energy, resulting in a plunge towards the nucleus. This heuristic argument can also be developed in general relativity in the widely-separated, slow-motion limit, with varying levels of rigor (see [25, 26] and references therein, and [27] for more recent developments up to 4.5 post-Newtonian order). Such flux-balance laws have provided a crucial test of general relativity from the measurement of the changing orbital period of the Hulse-Taylor binary pulsar [28].
@@
These examples highlight a key feature of flux-balance laws: they allow for the direct computation not of (even averaged) position or velocity, but of the slowly-evolving, no-longer-conserved quantities of the background motion. In the Kerr spacetime, for example, there are four constants of motion for a geodesic: the square of the mass $m^{2}$, the energy $E$, the azimuthal angular momentum $L_{z}$, and the “Carter constant” $K$ [29]. The Carter constant, which like $m^{2}$ is quadratic in the four-momentum, is related to the square of the orbital angular momentum $L^{2}$ in the limit where the primary black hole has no spin. These are sufficient to determine geodesic motion, but when the body has spin, in addition to generalizations of each of these constants [30, 31], there exists a fifth constant of motion, which we call the Rüdiger constant. [^1] Elsewhere in the literature (for example, [32, 33]), this constant is referred to as the linear Rüdiger invariant, in contrast to the generalization of the Carter constant mentioned above, which is often called the quadratic Rüdiger invariant. For simplicity, we instead use the term “Carter constant” to refer to both the usual constant of motion for non-spinning particle motion and its generalization to spinning-body motion. As such, “Rüdiger constant” in this paper refers only to the constant of motion which is linear in the four-momentum and spin of the body. $Y$ [30] For any sort of flux-balance law to be useful, it must therefore provide average rates of change of these quantities.
@@
Two general approaches have been used in the literature in order to derive flux-balance laws. The first is to use the properties of Green’s functions in the Kerr spacetime to derive relationships between the rates of change of quasi-conserved quantities to the amplitudes of solutions to the Teukolsky equation, which can be written as terms defined at the horizon and infinity. This approach was pioneered in [34] for computations of $E$ and $L_{z}$ without spin, which was then (at a much later date) extended to the linear-in-spin case in [35] (whose derivation was later corrected in [16]). In order to determine the full motion without spin, the evolution of the Carter constant $K$ was computed (once again in terms of asymptotic quantities) in [36, 37], which was later extended in [1] to the case of arbitrary constants of motion for geodesic motion in Kerr. The conserved quantities considered in [1] were the action variables in a Hamiltonian formulation of geodesic motion, which they then show can be used to compute the evolution of $K$.
@@
Recently, however, it was shown in [41] that, at least in the context of a scalar field theory and in the absence of spin (where the evolution of the Carter constant was derived in [42]), one can write the average evolution of the action variables in terms of fluxes of an appropriate conserved current through a local worldtube around the body. This presented a partial physical explanation for the results of [1]. Moreover, by naïvely guessing how to extend this flux-balance law all the way to the horizon and null infinity, a derivation was presented in [41] for a flux-balance law, for scalar fields, that was exactly analogous to the results in [1].
This paper begins where [41] ends, and provides an extension of this local flux-balance law beyond the limited scope of that paper. As such, the goals of this paper are threefold:
@@
$\left\langle\delta\dot{\Gamma}^{A}(\tau^{\prime})\right\rangle_{\tau^{\prime}}=\Pi^{AB}\left[\mathcal{F}_{B}+\nabla_{B}\langle\mathcal{H}(X^{\prime})\rangle_{\tau^{\prime}}\right],$ (1)
Without describing the notation explicitly, we will briefly describe this equation’s general features: first, the left-hand side represents a notion of an average rate of change of quantities defined on the worldline, and is defined in Eqs. (22) and (23). For the right-hand side, the first term is a flux of a conserved current through null infinity and the horizon that is defined in terms of metric perturbations in Eq. (121) below, and the second term is a derivative of an average “conservative” Hamiltonian (defined in Sec. IV.2) which generically appears, and is important in the case of resonances [1].
Unfortunately, because of the presence of the conservative Hamiltonian, this equation does not constitute a flux-balance law relating rates of change on the worldline to asymptotic fluxes of a conserved current, as this term still involves computing the regularized metric perturbation, from which the effective metric $\breve{g}_{ab}$ is constructed, on the worldline. Moreover, the left-hand side is not directly related to averaged rates of change of perturbed conserved quantities. As described in more detail in Sec. II.1 below, it is a covariantly-defined average of the rate of change of $\delta\Gamma^{A}$, the first-order deviation of the self-forced motion from the background trajectory $\Gamma$ through phase space. However, in order to make this average covariantly defined on phase space [like the remaining terms in Eq. (1)], one needs to include a bitensor introduced in [41] called the Hamilton propagator, which significantly complicates the interpretation in the case of spin, as we discuss in Sec. V.2 below.
@@
assuming that the background motion is non-resonant. Here $\vartheta^{\alpha}$ represents the angle variables, $A^{\beta}{}_{\alpha}$ describes the failure of the coordinates $(\vartheta^{\alpha},P_{\alpha})$ to be canonical, and $(\partial_{\vartheta^{\beta}})^{A}\mathcal{F}_{A}$ is the angle-variable component of $\mathcal{F}_{A}$. The free index of $\mathcal{F}_{A}$ arises from a derivative acting on the metric perturbation, and so this equation implies that it is the derivative of the metric perturbation with respect to the angle variables $\vartheta^{\alpha}$ (in a manner which we make more precise through the body of this paper) which is relevant to computing the evolution of the constants of motion $P_{\alpha}$.
Unlike Eq. (1), Eq. (2) is a true flux-balance law, as the right-hand side only involves the computation of the metric perturbation $h_{ab}$ at null infinity and the horizon. Moreover, the left-hand side is the average rate of change of the perturbed conserved quantity $\delta P_{\alpha}$, in the usual sense. In particular, this equation holds for non-spinning motion in the Kerr spacetime, and is equivalent to the results of [1]. However, in the case of spin, there no longer exists a full set of action-angle variables on the entire phase space, in a sense which we describe in more detail in Sec. V.2 below. As such, our results are much more limited: while we can still recover flux-balance laws for generalizations of $m^{2}$, $E$, $L_{z}$, and $K$, the evolution of the Rüdiger constant $Y$ is less well-behaved. As mentioned above, while flux-balance laws for $E$ and $L_{z}$ have already been known [35, 16], the flux-balance law for $K$ is new to this work, and arises due to a “miracle” that we describe in more detail in Sec. V.2 below. While the results of this paper are somewhat abstract, they provide a path towards explicit expressions for these flux-balance laws, as discussed in [43] using recent results in [44].
@@
Following [41], for tensor fields on the phase space manifold $\mathcal{M}$, we use capital Latin letters from the beginning of the alphabet ($A$, $B$, etc.) to denote abstract indices, and capital Hebrew letters from the beginning of the alphabet ($\aleph$, $\beth$, etc.) for coordinate indices. We will also use unhatted Greek letters for indices associated with groupings of four coordinates on phase space, such as the angle variables $\vartheta^{\alpha}$ or the conserved quantities $P_{\alpha}$. A difficulty arises then in denoting the components of tensors along these directions: as in Eq. (2) above, we resort to contracting tensors with (for example) the vector fields $(\partial_{\vartheta^{\alpha}})^{A}$ and $(\partial_{P_{\alpha}})^{A}$, which form a basis dual to the usual one-forms $(\mathrm{d}\vartheta^{\alpha})_{A}$ and $(\mathrm{d}P_{\alpha})_{A}$. Finally, for quantities appearing on the spacetime manifold $M$, we typically denote corresponding quantities on the phase space manifold $\mathcal{M}$ with capitalized versions of the same symbol; for example, the worldline $\gamma$ is determined by a trajectory $\Gamma$ through phase space.
@@
Finally, parentheses and square brackets will typically have their conventional meanings (that is, dependence as a function vs. dependence as a functional/coincidence limits [13]), to which we add the additional notation that curly braces will be used to indicate dependence as a local, linear functionals (from [47]). In certain situations, however, when brackets of the same type would be nested, such as in $f(g(x))$, we follow the usual convention of cycling through alternative brackets; in this example, we would use $f[g(x)]$, even if $f$ is not a functional of its argument. That this replacement of brackets has occurred should be clear from the surrounding context.
@@ II.1 General framework
We denote the common parameter of $\gamma$ and $\Gamma$ by $\tau$, as it will represent the proper time of the body, so that $\dot{\gamma}^{a}$ will be normalized.
@@
$\{f,\{g,h\}\}+\{g,\{h,f\}\}+\{h,\{f,g\}\}=0;$ (5)
this is equivalent to the statement that the Schouten(-Nijenhuis) bracket of $\Pi^{AB}$ with itself vanishes (see Theorem 10.6.2 of [48]). The equations of motion for $\gamma$ are then given by a Hamiltonian formulation if there exists a function $H$ on $\mathcal{M}$ that satisfies Hamilton’s equations:
@@
Now, we consider the more general case of a pseudo-Hamiltonian system. Here, the idea is that what one would typically think of as a Hamiltonian is not simply a function of a point on phase space $X$, but a function also of a trajectory through phase space $\bar{\Gamma}$ which is determined by a point $\bar{X}$ through which $\bar{\Gamma}$ passes. Following [49, 50, 51], we call such a function a pseudo-Hamiltonian. In the pseudo-Hamiltonian generalization of Hamilton’s equations, one takes a derivative with respect to $X$ at $\bar{X}$ fixed, and then takes $\bar{X}$ and $X$ to lie along the same trajectory $\Gamma$ in phase space. These systems allow for dissipation, unlike “true” Hamiltonian systems [52]. [^3] Note that, while they sound somewhat exotic, the idea of pseudo-Hamiltonian systems is far from new (see the discussion and references in [52]). For an example from the gravitational radiation literature, the Burke-Thorne potential for the leading-order post-Newtonian radiation-reaction force (see the discussion in Chapter 12 of [24]), $\Phi^{RR}(x)=\frac{1}{5}x_{i}x_{j}\frac{\mathrm{d}^{5}I^{\langle ij\rangle}}{\mathrm{d}t^{5}},$ can be thought of as the potential energy part of a pseudo-Hamiltonian: the quadrupole moment $I^{ij}$ depends on the trajectory of the object undergoing radiation reaction, which is kept fixed when $\Phi^{RR}$ is differentiated with respect to $x^{i}$ to find the force {to see this, compare Eqs. (12.198) and (12.204) of [24]}.
@@
for some value of $\tau$ along some trajectory $\Gamma$ in $\mathcal{M}$ that passes through $X$ and obeys Hamilton’s equations for the background Hamiltonian $H$. We then define
@@
that is, when the small parameter $\varepsilon=0$, the system becomes Hamiltonian for the background motion that determines $\Upsilon$ [^4] In order to describe self-force in the so-called “self-consistent formalism” [53, 15], one would need to instead consider the Hamiltonian flow map associated with the exact motion in the definition of the pseudo-Hamiltonian, which we would denote $\Upsilon_{\Delta}(\varepsilon)$. However, since we are computing the first-order self-force, the self-consistent formalism is not necessary, and so we just use the background Hamiltonian flow map.
@@
$\left.\delta\Gamma^{A}\nabla_{A}f\right|_{\Gamma(\tau)}\equiv\left.\frac{\partial f[\Gamma(\tau;\varepsilon)]}{\partial\varepsilon}\right|_{\varepsilon=0}.$ (12)
@@
As shown in Sec. II.B.2 of [41], the following evolution equation for $\delta\Gamma^{A}$ follows from Eq. (11), using the notation of the pseudo-Hamiltonian formalism (which was only implicit in [41]):
@@
That is, the Poisson bracket structure is fixed [^5] When we say “fixed” here, we mean that it does not explicitly depend on $\varepsilon$; any possible implicit dependence due to being evaluated at a point which itself depends on $\varepsilon$ is allowed, as can be seen in the derivation in [41]. as a function of $\varepsilon$. Note that this is a “gauge-dependent” statement: it only holds in a certain class of coordinates [for example, those in Eq. (43)], and if one performs an $\varepsilon$-dependent coordinate transformation (depending, for example, on the metric perturbation), then Eq. (14) will necessarily change. It is for this reason that we work in coordinates related to those in Eq. (43) by field-independent coordinate transformations.
@@
$\nabla_{A}(f\circ\Upsilon_{\tau^{\prime}-\tau})=\Upsilon^{A^{\prime}}{}_{A}\nabla_{A^{\prime}}f.$ (16)
In other words, the Hamilton propagator is the pushforward of the Hamiltonian flow map $\Upsilon_{\Delta}$, [^6] While the pushforward of a mapping from a manifold into itself is not discussed in terms of bitensors in typical textbooks on general relativity (such as, for example, Appendix C of [45]), it is not unprecedented: see for example Sec. 1.4 of [54]. and is a bitensor that is a one-form at $\Gamma(\tau)$ and a vector at $\Gamma(\tau+\Delta)$. [^7] Note that this definition is slightly different from that originally presented in [41]: there, $\Upsilon^{A^{\prime}}{}_{A}$ was defined such that it was a bitensor at any two points $X$ and $X^{\prime}$ along a given trajectory $\Gamma$, by using a function which gave the difference $\Delta$ in proper times between these two points. This allows one to take derivatives of $\Upsilon^{A^{\prime}}{}_{A}$ with respect to either of $X$ and $X^{\prime}$, entirely independently. However, as we discovered while writing the present paper, such a feature is not required for the derivation of flux-balance laws, and so we do not include this part of the definition.
@@
Finally, we note that, in both Eqs. (23) and (25), care must be taken that the quantity being averaged does, in fact, have a finite and well-defined average. As we will see below in Sec. V.1, this is not always the case, for particular components of these equations; however, in such cases, it will be true that both the left- and right-hand sides of Eq. (23) will be ill-defined.